use the given conditions to write an equation for the line in point-slope form and slope intercept form passing through (5,-9) and perpendicular to the line whose equation is x+2y=5
Answer by stanbon(75887)

use the given conditions to write an equation for the line in point-slope form and slope intercept form passing through (5,-9) and perpendicular to the line whose equation is x+2y=5
----------
The given line:
x + 2y = 5
2y = -x + 5
y = (-1/2)x + (5/2)
---
slope = -1/2
-------------------
Perpendicular line:
slope = 2
point: (5,-9)
----
Form: y = mx + b
Solve for "b":
-9 = 2*5 + b
b = -19
--------
Equation:
y = 2x - 19
=====================
Cheers,
Stan H.
====================